1. The CFO of MediSearch plc believes that investing into a project which will develop a new drug for fighting the flu virus, promises a long-term annual return of 8%. The expected return of the market index is RM = 17.9%, the volatility of the market index is σM  = 9.5%, and the risk-free asset earns Rf = 5.2%. The CFO has studied similar projects of other companies and believes that the covariance between the returns of this project and the returns of the market is σP,M  = 0.0104 (in decimal form). What is the beta of the project? Enter your answer rounded to two decimal places.

2.Lucas builds a portfolio by investing in two stocks only: Samsung (SMSN) and Huawei (SHE). According to the CAPM, the expected risk premium (i.e., the expected return minus the risk-free rate) of SMSN is 11.98% and the expected risk premium of SHE is 5.23%. The beta of SMSN is equal to 1.23. If Lucas puts 78% of his money in SMSN stock and 22% in SHE stock, what is the approximate beta of his portfolio?

3. Obi-wan forms an aggressive growth portfolio by investing 22% of his savings in Ford stock, 23% in Nissan stock, 21% in Kia stock, 13% in an index fund, and the last 21% is allocated on a bond fund. Assume for simplicity that the index fund is a good proxy to the market portfolio and has a beta equal to 1, whereas the bond fund is a good proxy to the riskless asset. The beta of Ford stock is 1.27, the beta of Nissan is 1.51, and the beta of Kia is 1.72. If the expected return of the market index is 14% and the risk-free asset yields 7%, what are the beta and the expected return of Obi-wan’s portfolio? Give your answer rounded to two decimal places.

What is the beta of the portfolio?
Answer #1

1) Beta of Project =Covariance of Project and return /(Standard Deviation of market) =0.0104/9.5% =0.11

2) Market Risk Premium =Expected Risk premium of SHE/Beta =5.23%/1.23=4.2520%
Beta of SMSN =Expected Risk Premium/Market Risk Premium =11.98%/4.252% =2.8175
Beta of Portfolio =78%*2.8175+22%*1.23 =2.47